### v2

@@ 3.6.2 Abbreviation expansion and self-correction module
Abbreviations are also a challenging part of molecular recognition. When describing molecular structures, abbreviations are commonly used to represent complex functional groups, such as ”Ph” for phenyl, ”Ac” for acetyl, and ”Bn” for benzyl. These abbreviations act as ”superatoms” in molecular diagrams.
To deduce the complete molecular structure from these abbreviated forms, additional processing steps are required. Current methodologies, whether rule-based or machine learning-driven, typically involve creating a list of common abbreviations along with their corresponding functional group structures and SMLIES, using this list to substitute the superatoms during model inference. However, due to the combinatorial nature of chemical abbreviations, relying solely on a list to cover all possible combinations is impractical. For instance, combinations such as ”BnO” or ”OBn” refer to the same structure but in a different order, or more random and longer abbreviations such as ”NHCOOH”, although considered hyperatomic, are almost impossible to enumerate exhaustively.
In practice, we also found that if the model makes minor mistakes in recognizing hyperatomic characters such as identifying ”OTBMS” as ”OTTBMS”, it can lead to incorrect prediction of the entire molecular structure.
We propose an abbreviation expansion and self-correction module, which is more generalized and robust than previous methods. Although We also have a list of common abbreviations and their corresponding SMILES, our approach can also be extended to unseen forms. In an abbreviation expansion process, we first 1) check whether the superatom is in the list, and if so, we directly output the corresponding SMILES; 2) otherwise, we split the superatom symbol into atomic characters, such as expanding ”O2CH3” into ”OOCHHH”. 3) After deriving the list of atoms, we greedily connect them based on valence bonds according to the SMILES formalism until their valence bonds are full, at which point the output is the final SMILES. 4) If SMILES cannot be successfully output, we will compare the original abbreviation symbol with every abbreviation in the list, find the abbreviation with the highest similarity, and replace it with the corresponding SMILES if the similarity is greater than the set threshold $\sigma=0.8$.
@@ 4.1.1 Train data
PubChem database [37] contains about 100M molecules. We use this database as the synthetic data. However, they are not all required, we estimate that we need about 1M structures. We randomly select 1M molecules from the database and render their images using chemical tools.
@@
Following previous work [3] which collected 680 K examples from USPTO [38] which contain molecular images and structure labels. We use this dataset as the real data. Coordinate annotations of the data are obtained by normalizing the relative coordinates available in MOLfiles according to the image size. This dataset is noisy and contains a variety of styles, which is closer to the actual use of the model than the synthetic data.
@@ 4.1.2 Test data
We evaluate our MolNexTR on five public realistic datasets, which are CLEF, UOB, USPTO, Staker, and ACS. The ACS dataset is a new dataset collected by [3] with 331 molecular images taken from ACS publications. Compared to other datasets, it is more diverse in terms of drawing styles and the use of abbreviations. Furthermore, we created three synthetic datasets, generated by indigo, RDKit, and ChemiDraw respectively. They were rendered from the same dataset of 5,719 molecules. Table 1 presents the details of the test sets we used.
Table 1: Summary of Test datasets. ✓ means inclusion, ✗ means no inclusion.
@@
CLEF realistic 992
UOB realistic 5,740
USPTO realistic 5,719
Staker realistic 50,000
ACS realistic 331
@@ 5 Results and Discussion
5.1 Comparison with State-of-the-Art Methods
Table 2: Result comparisons (%) of the overall SMILES sequence exact matching accuracy with the state-of-the-art methods on different test sets. “—” denotes that the results are unavailable.
@@ 5.1 Comparison with State-of-the-Art Methods
Indigo ChemDraw RDKit CLEF UOB USPTO Staker ACS
Rule-based MoIVec [40] 95.4 87.9 88.7 82.8 80.6 88.4 0.8 47.4
OSRA [17] 95.0 87.3 88.2 84.6 78.5 87.4 0.0 55.3
@@
Image2Graph [6] 51.7 82.9 55.1
SwinOCSR [5] 74.0 79.6 77.3 30.0 44.9 27.9 27.5
Img2Mol [7] 58.9 46.4 44.2 18.3 68.7 26.3 17.0 23.0
DECIMER [2] 69.6 86.1 82.3 62.7 88.2 41.1 40.8 46.5
MolScribe [3] 97.5 93.8 94.6 88.3 87.9 92.6 86.9 71.9
Ours 97.8 95.1 96.4 90.4 88.5 93.8 88.3 81.9
Table 3: Result comparisons (%) of the overall SMILES sequence exact matching accuracy with the state-of-the-art methods on different perturbed test sets. “—” denotes that the results are unavailable.
@@
CLEF UOB USPTO Staker ACS CLEF UOB USPTO Staker ACS
Rule-based MoIVec [40] 43.7 74.5 29.7 5.0 10.6 40.8 72.6 22.9 4.1 5.3
OSRA [17] 11.5 68.3 4.0 4.6 12.3
@@
Img2Mol [7] 21.1 74.9 29.7 51.7 8.6
DECIMER [2] 70.6 87.3 46.4 47.9 20.1 59.8 42.4
MolScribe [3] 90.4 86.7 92.5 65.0 53.1 87.1 81.6 88.9 73.0 62.8
Ours 90.4 88.4 93.2 78.9 68.8 89.5 84.3 91.2 73.2 70.8
Figure 5: Bar chart comparison of the SMILES sequence exact matching accuracy on molecules with chirality or abbreviations with MolScribe on different test sets.
@@
We compare our MolNexTR with the state-of-the-art (SOTA) methods including rule-based OSRA [17] and MolVec [40], and deep learning-based models Img2Mol [7], DECIMER [4], SwinOCSR [5], MSE-DUDL [1], ChemGrapher [28], Image2Graph [6], and MolScribe [3]. We directly use the results reported in previous works. The results are shown in Table 2. The rules-based systems MolVec and OSRA achieve good performance on CLEF, UOB, and USPTO, and decline on ACS due to the very diverse drawing styles of this dataset. However, they dropped significantly on Staker, possibly due to the very low resolution of this dataset. Among deep learning methods, we can observe that for synthetic datasets, MolNexTR achieves 97.8% in Indigo, 95.1% in ChemDraw and 96.4% in RDKit, which outperforms the SOTA methods by 0.3%, 1.3% and 1.8% respectively. On realistic datasets, our MolNexTR achieves 90.2% in CLEF, 88.5% in UOB, 93.8% in USPTO, and 88.3% in Staker respectively, which outperforms the SOTA methods by 1.1%, 0.6%, 1.2% and 1.4%, respectively, validating the superior performance of our innovative model architecture. Furthermore, our MolNexTR achieves 81.9% in ACS, which outperforms the SOTA methods by 10.0%, which is a very large improvement compared to the SOTA method. Compared with other datasets, the ACS has very diverse image styles and contains a lot of pollution. This proves that our series of data augmentation methods solve these problems well and improve the generalization and robustness of the model.
@@
We also evaluate the model on perturbed datasets with some image transform following the setup of Clevert et al., the results are shown on the left of Table 3. MolNexTR performs better than the SOTA methods and the accuracy decays less. We further construct another perturbed dataset with curved arrows on it to simulate the molecules in the mechanism images, which is also a molecular pattern commonly found in the real literature, the results are shown on the right of Table 3. MolNexTR can better recognize such molecular patterns without being affected by the arrows in the molecules though we haven’t included such molecules in the training process. These results further demonstrate that our model has excellent robustness to disturbances.
@@
We further compare the performance of MolNexTR and the SOTA method MolScribe on molecules with chirality and abbreviation in Fig. 5. MolNexTR shows better performance, especially on datasets with more diverse drawing styles. This is because we consider the presence of functional groups when judging chirality, and obtain the information of functional groups before judging chirality, which is not considered by MolScribe. When expanding abbreviations, the self-correction mechanism can improve the accuracy of judgment. On the other hand, our data augmentation method enables the model to have a better ability of optical character recognition (OCR).
@@
Table 4: Ablation study results on the superiority of dual-stream encoder. * means the component achieves significant performance improvement over 0.8% with p ¡ 0.05 via paired t-test.
@@
[…]
Table 5: Ablation study results on the effectiveness of data augmentation algorithm. * means the component achieves significant performance improvement over 2% with p $<$ 0.05 via paired t-test.
@@
[…]
@@ 5.2.2 Effectiveness of data augmentation algorithm
Figure 6: Visualizations on the ACS dataset compared to the SOTA method.
In this study, we conducted an experimental analysis to explore the effectiveness of each component in the data augmentation algorithm. The evaluated variants include $+$Render Augmentation, $+$Image Augmentation, $+$Molecular Augmentation, and $+$Image Contaminate Algorithm. Our results, presented in Table  5, demonstrate that each of these data augmentations can significantly improve the baseline performance. Specifically, when applied to the MolNexTR, $+$Render Aug. (Render Augmentation), $+$Image Aug. (Image Augmentation), $+$Mol Aug. (Molcular Augmentation), and $+$Img Contam Alg. (Image Contamination Algorithm) led to performance gains of 1.8%, 1.6%, 1.4% and 1.0% in Indigo, 4.7%, 3.5%, 5.8% and 1.5% in ChemDraw, 4.0%, 2.6%, 6.9% and 1.3% in CLEF, 4.2%, 3.9%, 6.2% and 6.7% in ACS, respectively. We can observe that for the in-domain dataset Indigo(which uses the same tools as the training data), the data augmentation method does not improve the accuracy much. However, once applied to the out-of-domain dataset, most components of the data augmentation can greatly improve the prediction accuracy, especially in the ACS dataset. Although the application of the image contamination algorithm has only little improvement on indigo, ChemDraw, and CLEF, it has great improvement on ACS datasets with more diverse styles and more pollution. It can be said that ACS data sets are more closely related to real-world situations, so it is still necessary to apply the image contaminate algorithm. The above demonstrates the effectiveness of our data augmentation methods. They improve the robustness and generalization of the model that can be better applied in the real world.
@@ 5.3 Qualitative Results
5.3.1 Visual Comparison with State-of-the-Art Methods
Figure 7: Visualizations of some hand-drawn molecules images compared to the SOTA method.
Figure 8: Visualizations of some molecules in mechanism images compared to the SOTA method.
In Fig. 6, we show some visualizations of the ACS dataset compared to the SOTA method MolScribe. The experimental results show that the proposed MolNexTR method can accurately predict molecular structure from a variety of molecular images with various styles and styles especially those with contamination or complex abbreviations compared to MolScribe. The results demonstrate the importance of our series of data augmentation algorithms and abbreviation expansion and self-correction modules.
@@ 5.3.2 Generalization of MolNexTR
To further verify the generalization of our model, we also compared the model with the SOTA method on some out-domain real hand-drawn molecules and molecules in mechanism images. Fig. 7 shows the visualizations of some hand-drawn molecules images compared to the SOTA method. Although the pattern of hand-drawn molecules is very different from the pattern of common molecular images (more random structures, more ambiguous atomic symbols, etc.), and we did not include any hand-drawn molecules in training, the experimental results show that our model performs better at recognizing all the bonds and atomic symbols compared to the SOTA method.
Fig. 8 shows the visualizations of some molecules in mechanism images compared to the SOTA method. MolNexTR can better recognize arrow symbols present in such molecules and is not affected by them, while the SOTA method often obfuscates additional arrow or atom symbols as the main part of the molecular and outputs wrong predictions. These results prove that MolNexTR has stronger generalization compared to the SOTA methods.
@@ 5.3.3 Error Cases Analysis
Figure 9: Visualization of some error cases.
We selected several cases with incorrect predictions for analysis. It is shown in Fig. 9. We can see that MolNexTR tends to make wrong predictions for extremely complex molecules. The main reasons are these molecules have some rare structures, and the number of pixels assigned to each atom is low compared to other molecules. There are also cases where the drawing style causes the model to fail to recognize chirality, such as when dashed and solid wedge are replaced by broken lines. In addition, for some more complex hand-drawn molecules and molecules in mechanism images, MolNexTR still cannot guarantee to output completely correct prediction results though we did not include such molecules in training. We hope that future work can continue to address these issues.
6 Conclusion

### v3

@@ 3.6.2 Abbreviation expansion and self-correction module
Abbreviations are also a challenging part of molecular recognition. When describing molecular structures, abbreviations are commonly used to represent complex functional groups, such as "Ph" for phenyl, "Ac" for acetyl, and "Bn" for benzyl. These abbreviations act as "superatoms" in molecular diagrams.
To deduce the complete molecular structure from these abbreviated forms, additional processing steps are required. Current methodologies, whether rule-based or machine learning-driven, typically involve creating a list of common abbreviations along with their corresponding functional group structures and SMILES, using this list to substitute the superatoms during model inference. However, due to the combinatorial nature of chemical abbreviations, relying solely on a list to cover all possible combinations is impractical. For instance, combinations such as "BnO" or "OBn" refer to the same structure but in a different order, or more random and longer abbreviations such as "NHCOOH", although considered hyperatomic, are almost impossible to enumerate exhaustively.
In practice, we also found that if the model makes minor mistakes in recognizing hyperatomic characters such as identifying "OTBMS" as "OTTBMS", it can lead to incorrect prediction of the entire molecular structure.
We propose an abbreviation expansion and self-correction module, which is more generalized and robust than previous methods. Although we also have a list of common abbreviations and their corresponding SMILES, our approach can also be extended to unseen forms. In an abbreviation expansion process, we first 1) check whether the superatom is in the list, and if so, we directly output the corresponding SMILES; 2) otherwise, we split the superatom symbol into atomic characters, such as expanding "O2CH3" into "OOCHHH". 3) After deriving the list of atoms, we greedily connect them based on valence bonds according to the SMILES formalism until their valence bonds are full, at which point the output is the final SMILES. 4) If SMILES cannot be successfully output, we will compare the original abbreviation symbol with every abbreviation in the list, find the abbreviation with the highest similarity, and replace it with the corresponding SMILES if the similarity is greater than the set threshold $\sigma=0.8$.
@@ 4.1.1 Train data
PubChem database [37] contains about 100M molecules. We use this database as the synthetic training data. However, they are not all required, we estimate that we need about 1M structures. We randomly select 1M molecules from the database and render their images using chemical tools.
@@
Following previous work [3], we collect 0.68M examples from USPTO [38] which contain molecular images and structure labels. We use this dataset as the real training data. Coordinate annotations of the data are obtained by normalizing the relative coordinates available in MOLfiles according to the image size. This dataset is noisy and contains a variety of styles, which is closer to the actual use of the model than the synthetic data.
@@ 4.1.2 Test data
We evaluate our MolNexTR on six public realistic datasets, which are CLEF, UOB, JPO, USPTO, Staker, and ACS. The ACS dataset is a new dataset collected by [3] with 331 molecular images taken from ACS publications. Compared to other datasets, it is more diverse in terms of drawing styles and the use of abbreviations. Furthermore, we created three synthetic datasets, generated by indigo, RDKit, and ChemiDraw, respectively. They were rendered from the same dataset of 5,719 molecules. Table 1 presents the details of the test sets we used.
Table 1: Summary of the test datasets
@@
CLEF real 992
UOB real 5,740
JPO real 450
USPTO real 5,719
Staker real 50,000
ACS real 331
00footnotetext: ✓ means inclusion, ✗ means no inclusion.
@@ 5 Results and Discussion
5.1 Comparison with Current Methods
Table 2: Comparison of our model’s results with current methods across various test sets
@@ 5.1 Comparison with Current Methods
Indigo ChemDraw RDKit CLEF UOB JPO USPTO Staker ACS
Rule-based MoIVec [40] 95.4 87.9 88.7 82.8 80.6 67.8 88.4 0.8 47.4
OSRA [18] 95.0 87.3 88.2 84.6 78.5 55.3 87.4 0.0 55.3
@@
Image2Graph [6] 51.7 82.9 50.3 55.1
SwinOCSR [5] 74.0 79.6 77.3 30.0 44.9 13.8 27.9 27.5
Img2Mol [7] 58.9 46.4 44.2 18.3 68.7 16.4 26.3 17.0 23.0
DECIMER [2] 69.6 86.1 82.3 62.7 88.2 55.2 41.1 40.8 46.5
MolScribe [3] 97.5 93.8 94.6 88.3 87.9 77.7 92.6 86.9 71.9
Ours 97.8 95.1 96.4 90.4 88.5 82.1 93.8 88.3 81.9
00footnotetext: Scores are in overall SMILES sequence exact matching accuracy (%). "—" denotes that the results are unavailable. Bold represents the best performance and underline represents the second-best performance
Table 3: Comparison of our model’s results with current methods across various perturbed test sets
@@
CLEF UOB JPO USPTO Staker ACS CLEF UOB JPO USPTO Staker ACS
Rule-based MoIVec [40] 43.7 74.5 22.5 29.7 5.0 10.6 40.8 72.6 20.6 22.9 4.1 5.3
OSRA [18] 11.5 68.3 10.8 4.0 4.6 12.3
@@
Img2Mol [7] 21.1 74.9 8.9 29.7 51.7 8.6
DECIMER [2] 70.6 87.3 34.1 46.4 47.9 20.1 59.8 38.5 42.4
MolScribe [3] 90.4 86.7 52.0 92.5 65.0 53.1 87.1 81.6 56.7 88.9 73.0 62.8
Ours 90.4 88.4 71.9 93.2 78.9 68.8 89.5 84.3 67.7 91.2 73.2 70.8
00footnotetext: Scores are in overall SMILES sequence exact matching accuracy (%). "—" denotes that the results are unavailable. Bold represents the best performance and underline represents the second-best performance
Figure 5: Comparison of our model’s results with MolScribe on molecules with chirality or abbreviations.
@@
We compare our MolNexTR with the current methods including rule-based OSRA (Version 2.1.3) [18] and MolVec (Version 0.9.8) [40], and deep learning-based models Img2Mol [7], DECIMER (Version 2.1.0) [4], SwinOCSR [5], MSE-DUDL [1], ChemGrapher [29], Image2Graph [6], and MolScribe (The best checkpoint) [3]. We directly use the results reported in previous works. The results are shown in Table 2. The rules-based systems MolVec and OSRA achieve good performance on CLEF, UOB, and USPTO, but decline on JPO due to the Japanese characters involved. The performance on ACS decreases more due to the extremely diverse drawing styles of this dataset. However, they dropped significantly on Staker, possibly due to the low resolution of this dataset. Among deep learning methods, we can observe that for synthetic datasets, MolNexTR achieves 97.8% in Indigo, 95.1% in ChemDraw and 96.4% in RDKit, which outperforms the second-best method by 0.3%, 1.3% and 1.8% respectively. On realistic datasets, our MolNexTR achieves 90.2% in CLEF, 88.5% in UOB, 82.1% in JPO, 93.8% in USPTO, and 88.3% in Staker, respectively, which outperforms the second-best method by 1.1%, 0.3%, 4.4%, 1.2% and 1.4%, respectively, validating the superior performance of our innovative model architecture. Notably, our MolNexTR achieves 81.9% in ACS, which outperforms the second-best method by 10.0%, which is a significant improvement. Compared with other datasets, the ACS has more diverse image styles and contains much more contamination. This proves that our series of data augmentation methods solve these problems well and improve the generalization and robustness of the model.
@@
We also evaluate the model on perturbed datasets with some image transform following the setup of Clevert et al., the results are shown on the left of Table 3. MolNexTR performs better than the current methods, and the accuracy decays less. We further construct another perturbed dataset with curved arrows on it to simulate the molecules in the mechanism images, which is also a molecular pattern commonly found in the real literature, the results are shown on the right of Table 3. MolNexTR can better recognize such molecular patterns without being affected by the arrows in the molecules though we haven’t included such molecules in the training process. These results further demonstrate that our model has excellent robustness to deal with disturbances.
@@
We further compare the performance on molecules with chirality and abbreviation between MolNexTR and the MolScribe, which performs the best on multiple test datasets on average among current methods. The results are shown in Fig. 5. MolNexTR shows better performance, especially on datasets with more diverse drawing styles. This is because we consider the presence of functional groups when judging chirality, and obtain the information of functional groups before judging chirality, which is not considered by MolScribe. When expanding abbreviations, the self-correction mechanism can improve the accuracy of judgment. On the other hand, our data augmentation method enables the model to have a better OCR ability.
@@
Table 4: Ablation study results on the superiority of the dual-stream encoder
@@
97.0+0.3 94.5+0.7 89.4+0.6 80.8+0.5
97.8${}^{*}_{{+0.8}}$ 95.1+0.6 90.2${}^{*}_{{+0.8}}$ 81.9${}^{*}_{{+1.1}}$
00footnotetext: ✓ means inclusion, ✗ means no inclusion. * means the component achieves significant performance improvement over 0.8% with p $<$ 0.05 via paired t-test.
Table 5: Ablation study results on the effectiveness of the data augmentation algorithm
@@
93.8+1.8 85.3${}^{*}_{+4.7}$ 79.4${}^{*}_{+4.0}$ 65.0${}^{*}_{+4.2}$
95.4+1.6 88.8${}^{*}_{+3.5}$ 82.0${}^{*}_{+2.6}$ 68.9${}^{*}_{+3.9}$
96.8+1.4 93.6${}^{*}_{+5.8}$ 88.9${}^{*}_{+6.9}$ 75.1${}^{*}_{+6.2}$
97.8+1.0 95.1+1.5 90.2+1.3 81.9${}^{*}_{+6.7}$
00footnotetext: ✓ means inclusion, ✗ means no inclusion. * means the component achieves significant performance improvement over 2% with p $<$ 0.05 via paired t-test.
@@ 5.2.2 Effectiveness of data augmentation algorithm
Figure 6: Visualization of our model on the ACS dataset compared to MolScribe.
In this study, we conducted an experimental analysis to explore the effectiveness of each component in the data augmentation algorithm. The evaluated variants include $+$Render Augmentation, $+$Image Augmentation, $+$Molecular Augmentation, and $+$Image Contaminate Algorithm. Our results, presented in Table  5, demonstrate that each of these data augmentations can significantly improve the baseline performance. Specifically, when applied to the MolNexTR, $+$Render Aug. (Render Augmentation), $+$Image Aug. (Image Augmentation), $+$Mol Aug. (Molcular Augmentation), and $+$Img Contam Alg. (Image Contamination Algorithm) led to performance gains of 1.8%, 1.6%, 1.4% and 1.0% in Indigo, 4.7%, 3.5%, 5.8% and 1.5% in ChemDraw, 4.0%, 2.6%, 6.9% and 1.3% in CLEF, 4.2%, 3.9%, 6.2% and 6.7% in ACS, respectively. We can observe that for the in-domain dataset Indigo(which uses the same tools as the training data), the data augmentation method does not improve the accuracy much. However, once applied to the out-of-domain dataset, most components of the data augmentation can greatly improve the prediction accuracy, especially in the ACS dataset. Although the application of the image contamination algorithm has only slight improvement on indigo, ChemDraw, and CLEF, it has great improvement on ACS datasets with more diverse styles and more pollution. It can be said that ACS data sets are more closely related to real-world situations, so it is still necessary to apply the image contaminate algorithm. The above demonstrates the effectiveness of our data augmentation methods. They improve the robustness and generalization of the model that can be better applied in the real world.
@@ 5.3 Qualitative Results
5.3.1 Visual Comparison with Current Methods
Figure 7: Visualization of our model on some hand-drawn molecule images compared to MolScribe.
Figure 8: Visualization of our model on some molecules in mechanism images compared to MolScribe.
In Fig. 6, we show some visualizations of the ACS dataset compared to the second-best method MolScribe. The experimental results show that the proposed MolNexTR method can accurately predict molecular structure from a variety of molecular images with various styles, especially those with contamination or complex abbreviations compared to MolScribe. The results demonstrate the importance of our series of data augmentation algorithms and abbreviation expansion and self-correction modules.
@@ 5.3.2 Generalization of MolNexTR
To further verify the generalization of our model, we also compared the model with the second-best method on some out-domain real hand-drawn molecules and molecules in mechanism images. Fig. 7 shows the visualizations of some hand-drawn molecule images compared to the second-best method. Although the pattern of hand-drawn molecules is very different from the pattern of common molecular images (more random structures, more ambiguous atomic symbols, etc.), and we did not include any hand-drawn molecules in training, the experimental results show that our model performs better at recognizing all the bonds and atomic symbols compared to the current methods.
Fig. 8 shows the visualizations of some molecules in mechanism images compared to the second-best method. MolNexTR can better recognize arrow symbols present in such molecules and is not affected by them, while the other method often obfuscates additional arrow or atom symbols as the main part of the molecular and outputs wrong predictions. These results prove that MolNexTR has stronger generalization compared to the current methods.
@@ 5.3.3 Error Cases Analysis
Figure 9: Visualization of our model on some error cases.
We selected several cases with incorrect predictions for analysis. It is shown in Fig. 9. We can see that MolNexTR tends to make wrong predictions for extremely complex molecules. The main reasons are these molecules have some rare structures, and the number of pixels assigned to each atom is low compared to other molecules. There are also cases where the drawing style causes the model to fail to recognize chirality, such as when dashed and solid wedges are replaced by broken lines. In addition, for some more complex hand-drawn molecules and molecules in mechanism images, MolNexTR still cannot guarantee the output of completely correct prediction results, though we did not include such molecules in training. We hope that future work can continue to address these issues.
6 Conclusions